# Learning on One Mode: Addressing Multi-Modality in Offline Reinforcement Learning

### v1

@@ 1 Introduction
Offline reinforcement learning (RL) enables policy learning from static datasets, without active environment interaction, making it ideal for high-stakes applications like autonomous driving and robot manipulation [Levine et al., 2020, Ma et al., 2022, Wang et al., 2024a]. A key challenge in offline RL is managing the discrepancy between the learned policy and the behaviour policy that generated the dataset. Small discrepancies can hinder policy improvement, while large discrepancies push the learned policy into uncharted areas, causing significant extrapolation errors and poor generalisation [Fujimoto et al., 2019, Yang et al., 2023]. Addressing these challenges, existing research has proposed various solutions. Conservative approaches penalise actions that stray into out-of-distribution (OOD) regions [Yu et al., 2020, Kumar et al., 2020], while others regularise the policy by minimising its divergence from the behaviour policy, ensuring better fidelity to the dataset [Fujimoto and Gu, 2021, Wu et al., 2019]. Another solution is weighted imitation learning, which aims to replicate actions from the dataset through supervised learning techniques [Mao et al., 2023, Peng et al., 2019].
@@
Many real-world datasets introduce an additional challenge: multi-modal action distributions [Wang et al., 2024b, Chen et al., 2022, Zhou et al., 2020]. These datasets are common in practice because they often integrate data from diverse sources, such as multiple policies, human demonstrations, or distinct exploration strategies. This diversity arises naturally in domains where the same task can be approached in various ways, leading to states with multiple valid but potentially conflicting actions. For instance, in autonomous driving, different driving styles — conservative versus aggressive — may lead to different but equally valid ways of navigating a road. Similarly, in robotic manipulation, an object can be grasped in various ways depending on the robot’s approach, the object’s position, and environmental constraints. This multi-modality is not an exception but rather a frequent occurrence in complex, real-world decision-making tasks, as systems often integrate experience from various sources to handle different scenarios. Thus, it becomes crucial to model and manage these multi-modal action distributions effectively.
@@
Most offline RL approaches implicitly assume a unimodal action distributions, which can force policies to converge toward an average action that may not exist in the dataset, leading to degraded performance. This limitation is particularly evident in scenarios where policies fail to capture complex multi-modal distributions, instead collapsing into suboptimal or invalid averages, as studied by Cai et al. [2023], Wang et al. [2024b], Yang et al. [2022b] and illustrated in Appendix A. Recent work has attempted to address this by modelling the full multi-modal action distribution using expressive generative models such as GANs, VAEs, and diffusion models [Zhou et al., 2020, Chen et al., 2022, Wang et al., 2023]. However, these models often overcomplicate the learning process by capturing the entire action distribution, which is unnecessary when only a subset of the modes is relevant for optimal decision-making.
@@
We propose a simpler yet effective approach: Weighted Imitation Learning on One Mode (LOM). Our insight is that learning from a single mode — the one with the highest expected return — is sufficient to generate optimal actions. Instead of modeling the full multi-modal distribution, LOM identifies and focuses on the most promising mode for each state.
@@
Figure 1: The three steps of LOM. (1) Learn a network producing the parameters of a GMM to model the behaviour policy. (2) Evaluate each mode via the expected return of its actions and the select the optimal mode $\phi^{1}$. (3) Sample actions from $\phi^{1}$ for weighted imitation learning.
@@ 2 Related Work
The primary challenge in offline RL lies in managing the distribution shift between the behaviour policy, which generated the offline dataset, and the learned policy. This shift can cause the learned policy to produce actions that are not well-represented in the offline dataset, leading to inaccurate value function estimations and degraded performance [Levine et al., 2020, Fujimoto et al., 2019]. To mitigate the risks associated with OOD actions, one approach involves using conservative value functions, which penalise actions that deviate from the behaviour policy [Fujimoto et al., 2019, Kumar et al., 2020, Bai et al., 2022, Sun et al., 2022, An et al., 2021]. Another strategy focuses on regularising the learned policy to ensure its proximity to the behaviour policy. This regularisation can be measured using metrics such as mean squared error (MSE) [Beeson and Montana, 2024, Fujimoto and Gu, 2021] or more sophisticated metrics like the Wasserstein distance, Jensen-Shannon divergence [Yang et al., 2022b, Wu et al., 2019], and weighted MSE [Ma et al., 2024]. These methods aim to keep the learned policy within a safer, well-understood operational space, reducing the likelihood of selecting OOD actions. For a comprehensive review of value penalties and policy regularisation techniques, we refer readers to [Wu et al., 2019]. In contrast to these approaches, LOM tackles the distribution shift problem by focusing on a single, optimal mode of the behaviour policy, thereby reducing the risk of OOD actions while still allowing for significant policy improvement.
@@
Our proposed method contributes to the growing field of weighted imitation learning, which revolves around two fundamental questions: which actions should be imitated and how should these actions be weighted [Brandfonbrener et al., 2021]. Most existing methods imitate actions directly from the dataset [Peng et al., 2019, Wang et al., 2018, Nair et al., 2021, Wang et al., 2020], while some approaches [Siegel et al., 2020, Mao et al., 2023] suggest imitating actions from the most recent policies. These methods typically weight actions based on their advantage, conditioned either on the behaviour policy [Peng et al., 2019, Wang et al., 2018] or on the most recently updated policy [Nair et al., 2021, Wang et al., 2020, Siegel et al., 2020, Mao et al., 2023]. However, a key limitation of these approaches is their underlying assumption that the imitated actions are drawn from a uni-modal Gaussian distribution, an oversimplification that often fails in complex, real-world datasets where multi-modal distributions are common [Lynch et al., 2020, Yang et al., 2022b]. LOM addresses this limitation by forgoing the uni-modal assumption, instead extracting a highly-rewarded Gaussian mode from the behaviour policy.
@@
Offline RL datasets are often collected from multiple, unknown policies, leading to states with several valid action labels, which may conflict with one another. Most existing methods employ a uni-modal Gaussian policy, which inadequately captures the inherent multi-modality in such datasets. To address this limitation, PLAS [Zhou et al., 2020] decodes variables sampled from a VAE latent space, while LAPO [Chen et al., 2022] leverages advantage-weighted latent spaces for further policy optimisation. Additionally, Yang et al. [2022b] demonstrate that GANs can model multiple action modes, and GOPlan [Wang et al., 2024b] extends this by applying exponentiated advantage weighting to highlight highly-rewarded modes. DAWOG [Wang et al., 2024a] and DMPO [Osa et al., 2023] separate the state-action space and learn a conditioned Gaussian policy or a mixture of deterministic policies to solve the regions individually. Techniques such as normalising flows and diffusion models can also model multi-modal distributions by progressively transforming an initial distribution into the target distribution [Wang et al., 2023, Akimov et al., 2022]. However, these techniques are computationally intensive, making their inference processes inefficient. Beyond direct policy improvements, approaches like TD3+RKL [Cai et al., 2023] exploit the mode-seeking property of reverse KL-divergence to narrow action coverage, while BAW [Yang et al., 2022a] filters actions with values in the leftmost quantile. Despite these advances, most methods attempt to capture or model the entire multi-modal distribution, leading to increased complexity. In contrast, LOM simplifies the learning process by extracting and focusing on a single, optimal mode from the behaviour policy, effectively addressing the multi-modality challenge without the need to model the entire distribution.
@@ 3 Preliminaries
The value function is defined as $V^{\pi}(s_{t})=\mathbb{E}_{a_{t}\sim\pi(\cdot\mid s_{t})}Q^{\pi}(s_{t},a_{t})$, and the advantage function is $A^{\pi}(s_{t},a_{t})=Q^{\pi}(s_{t},a_{t})-V^{\pi}(s_{t})$.
In offline RL, the agent learns exclusively from a fixed dataset $\mathcal{D}$ collected by one or more behaviour policies, denoted as $\pi_{b}$, without further interaction with the environment [Levine et al., 2020]. Our approach is based on weighted imitation learning, where the objective is to optimize
$J(\pi)=\arg\max_{\pi}\mathbb{E}_{s\sim d_{\pi_{b}},a\sim\pi_{b}(\cdot\mid s)}\left[\exp\left(\frac{1}{\beta}A^{\pi_{b}}(s,a)\right)\log\pi(a\mid s)\right],$
where $\beta$ is a hyper-parameter. This formulation encourages learning a policy $\pi$ that imitates actions from $\mathcal{D}$, weighted by the exponentiated advantage of the behaviour policy $\pi_{b}$, thereby implicitly constraining the KL-divergence between $\pi$ and $\pi_{b}$ [Wang et al., 2018]. Recent works have extended this approach by imitating actions from the current learned policy [Siegel et al., 2020, Mao et al., 2023], or weighting actions based on the advantage conditioned on the current policy [Nair et al., 2021, Wang et al., 2020].
@@ 4.1 Multi-modal Behavioural Policy
To address the challenge of multi-modal action distributions in offline RL, we model the behaviour policy as a Gaussian Mixture Model (GMM). GMMs are flexible in representing complex distributions and can approximate any density function with sufficient components [McLachlan and Basford, 1988], making them well-suited for capturing the heterogeneous nature of offline datasets that often arise from multiple policies or exploration strategies. The behaviour policy $\pi_{b}(a\mid s)$ is expressed as:
@@ 4.2 Hyper-Markov Decision Process
$\mathcal{P}_{H}(s_{t+1}\mid s_{t},u_{t})=\mathbb{E}_{a_{t}\sim\phi^{u_{t}}(a\mid s_{t})}\mathcal{P}(s_{t+1}\mid s_{t},a_{t}),$
$r_{H}(s_{t},u_{t})=\mathbb{E}_{a_{t}\sim\phi^{u_{t}}(a\mid s_{t})}r(s_{t},a_{t}).$
@@ 4.3 Hyper Q-function
The proof can be found in the Appendix B.1.
@@ 4.5 Weighted Imitation Learning on One Mode
Based on the greedy hyper-policy $\zeta_{g}$, we obtain the corresponding greedy policy $\pi_{\zeta_{g}}$, which, according to Theorem 1, is not worse than $\pi_{b}$. However, rather than simply imitating $\pi_{\zeta_{g}}$, we apply weighted imitation learning to further improve the policy. Specifically, we aim to find a policy $\pi$ that maximises the expected improvement $\eta(\pi)=J(\pi)-J(\pi_{\zeta_{g}})$. The expected improvement $\eta(\pi)$ can be expressed in terms of the advantage $A^{\pi_{\zeta_{g}}}(s,a)$ [Kakade and Langford, 2002, Schulman et al., 2015]: $\eta(\pi)=\mathbb{E}_{s\sim d_{\pi}(\cdot),a\sim\pi(\cdot\mid s)}[A^{\pi_{\zeta_{g}}}(s,a)]$. However, in offline RL, state samples from the distribution $d_{\pi}(s)$ are unavailable. Instead, we approximate the expected improvement by using an estimate $\hat{\eta}$ [Kakade and Langford, 2002]: $\hat{\eta}(\pi)=\mathbb{E}_{s\sim d_{\pi_{\zeta_{g}}}(\cdot),a\sim\pi(\cdot\mid s)}[A^{\pi_{\zeta_{g}}}(s,a)]$. This approximation provides a good estimate of $\eta(\pi)$ if $\pi$ and $\pi_{\zeta_{g}}$ are close in terms of KL-divergence [Schulman et al., 2015]. As a result, we can formulate the following constrained policy optimisation problem:
$\begin{split}&J(\pi)=\arg\max_{\pi}\sum_{s}d_{\pi_{\zeta_{g}}}(s)\sum_{a}\pi(a\mid s)A^{\pi_{\zeta_{g}}}(s,a),\\ &\text{s.t. }\sum_{s}d_{\pi_{\zeta_{g}}}(s)D_{KL}(\pi(\cdot\mid s)||\pi_{\zeta_{g}}(\cdot\mid s))\leq\epsilon.\end{split}$ (9)
@@
$J(\pi)$ $=\arg\min_{\pi}\mathbb{E}_{s\sim d_{\pi_{b}}(\cdot)}[D_{KL}(\pi^{*}(\cdot\mid s)||\pi(\cdot\mid s))]$ (10)
@@
In this framework, the actions to be imitated are sampled from $\pi_{\zeta_{g}}$, a uni-modal Gaussian policy, thereby addressing the multi-modality issue. Furthermore, actions within this mode are weighted according to their advantage, encouraging the policy to concentrate on the highly-rewarded actions. As shown in Theorem 2, the policy learned by LOM has a theoretical advantage over both $\pi_{\zeta_{g}}$ and the original behaviour policy $\pi_{b}$.
@@
This theorem implies that our method introduces a novel two-step improvement process. LOM achieves further policy improvement compared to one-step algorithms while requiring less off-policy evaluation than multi-step algorithms [Brandfonbrener et al., 2021].

### v2

@@ 1 Introduction
Offline reinforcement learning (RL) enables policy learning from static datasets, without active environment interaction, making it ideal for high-stakes applications like autonomous driving and robot manipulation (Levine et al., 2020; Ma et al., 2022; Wang et al., 2024a). A key challenge in offline RL is managing the discrepancy between the learned policy and the behaviour policy that generated the dataset. Small discrepancies can hinder policy improvement, while large discrepancies push the learned policy into uncharted areas, causing significant extrapolation errors and poor generalisation (Fujimoto et al., 2019; Yang et al., 2023). Addressing these challenges, existing research has proposed various solutions. Conservative approaches penalise actions that stray into out-of-distribution (OOD) regions (Yu et al., 2020; Kumar et al., 2020), while others regularise the policy by minimising its divergence from the behaviour policy, ensuring better fidelity to the dataset (Fujimoto & Gu, 2021; Wu et al., 2019). Another solution is weighted imitation learning, which aims to replicate actions from the dataset through supervised learning techniques (Mao et al., 2023; Peng et al., 2019).
@@
Many real-world datasets introduce an additional challenge: multi-modal action distributions (Wang et al., 2024b; Chen et al., 2022; Zhou et al., 2020). These datasets are common in practice because they often integrate data from diverse sources, such as multiple policies, human demonstrations, or distinct exploration strategies. This diversity arises naturally in domains where the same task can be approached in various ways, leading to states with multiple valid but potentially conflicting actions. For instance, in autonomous driving, different driving styles — conservative versus aggressive — may lead to different but equally valid ways of navigating a road. Similarly, in robotic manipulation, an object can be grasped in various ways depending on the robot’s approach, the object’s position, and environmental constraints. This multi-modality is not an exception but rather a frequent occurrence in complex, real-world decision-making tasks, as systems often integrate experience from various sources to handle different scenarios. Thus, it becomes crucial to model and manage these multi-modal action distributions effectively.
@@
Most offline RL approaches implicitly assume a unimodal action distribution, which can force policies to converge toward an average action that may not exist in the dataset, leading to degraded performance. This limitation is particularly evident in scenarios where policies fail to capture complex multi-modal distributions, instead collapsing into suboptimal or invalid averages, as studied byCai et al. (2023); Wang et al. (2024b); Yang et al. (2022b) and illustrated in Appendix A. Recent work has attempted to address this by modelling the full multi-modal action distribution using expressive generative models such as GANs, VAEs, and diffusion models (Zhou et al., 2020; Chen et al., 2022; Wang et al., 2023). However, these models often overcomplicate the learning process by capturing the entire action distribution, which is unnecessary when only a subset of the modes is relevant for optimal decision-making.
@@
We propose a simpler yet effective approach: Weighted Imitation Learning on One Mode (LOM). Our insight is that learning from a single mode — the one with the highest expected return — is sufficient to generate optimal actions. Instead of modelling the full multi-modal distribution, LOM identifies and focuses on the most promising mode for each state.
@@
Figure 1: The three steps of LOM. (1) Learn a network producing the parameters of a GMM to model the behaviour policy. (2) Evaluate each mode via the expected return of its actions and then select the optimal mode $\phi^{1}$. (3) Sample actions from $\phi^{1}$ for weighted imitation learning.
@@ 2 Related Work
The primary challenge in offline RL lies in managing the distribution shift between the behaviour policy, which generated the offline dataset, and the learned policy. This shift can cause the learned policy to produce actions that are not well-represented in the offline dataset, leading to inaccurate value function estimations and degraded performance (Levine et al., 2020; Fujimoto et al., 2019). To mitigate the risks associated with OOD actions, one approach involves using conservative value functions, which penalise actions that deviate from the behaviour policy (Fujimoto et al., 2019; Kumar et al., 2020; Bai et al., 2022; Sun et al., 2022; An et al., 2021). Another strategy focuses on regularising the learned policy to ensure its proximity to the behaviour policy. This regularisation can be measured using metrics such as mean squared error (MSE) (Beeson & Montana, 2024; Fujimoto & Gu, 2021) or more sophisticated metrics like the Wasserstein distance, Jensen-Shannon divergence (Yang et al., 2022b; Wu et al., 2019), and weighted MSE (Ma et al., 2024). These methods aim to keep the learned policy within a safer, well-understood operational space, reducing the likelihood of selecting OOD actions. For a comprehensive review of value penalties and policy regularisation techniques, we refer readers to (Wu et al., 2019). In contrast to these approaches, LOM tackles the distribution shift problem by focusing on a single, optimal mode of the behaviour policy, thereby reducing the risk of OOD actions while still allowing for significant policy improvement.
@@
Our proposed method contributes to the growing field of weighted imitation learning, which revolves around two fundamental questions: which actions should be imitated and how should these actions be weighted (Brandfonbrener et al., 2021). Most existing methods imitate actions directly from the dataset (Peng et al., 2019; Wang et al., 2018; Nair et al., 2021; Wang et al., 2020), while some approaches (Siegel et al., 2020; Mao et al., 2023) suggest imitating actions from the most recent policies. These methods typically weigh actions based on their advantage, conditioned either on the behaviour policy (Peng et al., 2019; Wang et al., 2018) or on the most recently updated policy (Nair et al., 2021; Wang et al., 2020; Siegel et al., 2020; Mao et al., 2023). However, a key limitation of these approaches is their underlying assumption that the imitated actions are drawn from a unimodal Gaussian distribution, an oversimplification that often fails in complex, real-world datasets where multi-modal distributions are common (Lynch et al., 2020; Yang et al., 2022b). LOM addresses this limitation by forgoing the unimodal assumption, instead extracting a highly-rewarded Gaussian mode from the behaviour policy.
@@
Offline RL datasets are often collected from multiple, unknown policies, leading to states with several valid action labels, which may conflict with one another. Most existing methods employ a unimodal Gaussian policy, which inadequately captures the inherent multi-modality in such datasets. To address this limitation, PLAS (Zhou et al., 2020) decodes variables sampled from a VAE latent space, while LAPO (Chen et al., 2022) leverages advantage-weighted latent spaces for further policy optimisation. Additionally, Yang et al. (2022b) demonstrate that GANs can model multiple action modes, and GOPlan (Wang et al., 2024b) extends this by applying exponentiated advantage weighting to highlight highly-rewarded modes. DAWOG (Wang et al., 2024a) and DMPO (Osa et al., 2023) separate the state-action space and learn a conditioned Gaussian policy or a mixture of deterministic policies to solve the regions individually. Techniques such as normalising flows and diffusion models can also model multi-modal distributions by progressively transforming an initial distribution into the target distribution (Wang et al., 2023; Akimov et al., 2022). However, these techniques are computationally intensive, making their inference processes inefficient. Beyond direct policy improvements, approaches like TD3+RKL (Cai et al., 2023) exploit the mode-seeking property of reverse KL-divergence to narrow action coverage, while BAW (Yang et al., 2022a) filters actions with values in the leftmost quantile. Despite these advances, most methods attempt to capture or model the entire multi-modal distribution, leading to increased complexity. In contrast, LOM simplifies the learning process by extracting and focusing on a single, optimal mode from the behaviour policy, effectively addressing the multi-modality challenge without the need to model the entire distribution.
@@ 3 Preliminaries
The value function is defined as $V^{\pi}(s_{t})=\mathbb{E}_{a_{t}\sim\pi(\cdot\mid s_{t})}[Q^{\pi}(s_{t},a_{t})]$, and the advantage function is $A^{\pi}(s_{t},a_{t})=Q^{\pi}(s_{t},a_{t})-V^{\pi}(s_{t})$.
In offline RL, the agent learns exclusively from a fixed dataset $\mathcal{D}$ collected by one or more behaviour policies, denoted as $\pi_{b}$, without further interaction with the environment (Levine et al., 2020). Our approach is based on weighted imitation learning, where the objective is to optimize
$J(\pi)=\mathbb{E}_{s\sim d_{\pi_{b}},a\sim\pi_{b}(\cdot\mid s)}\left[\exp\left(\frac{1}{\beta}A^{\pi_{b}}(s,a)\right)\log\pi(a\mid s)\right],$
where $\beta$ is a hyper-parameter. This formulation encourages learning a policy $\pi$ that imitates actions from $\mathcal{D}$, weighted by the exponentiated advantage of the behaviour policy $\pi_{b}$, thereby implicitly constraining the KL-divergence between $\pi$ and $\pi_{b}$ (Wang et al., 2018). Recent works have extended this approach by imitating actions from the current learned policy (Siegel et al., 2020; Mao et al., 2023), or weighting actions based on the advantage conditioned on the current policy (Nair et al., 2021; Wang et al., 2020).
@@ 4.1 Multi-modal Behavioural Policy
To address the challenge of multi-modal action distributions in offline RL, we model the behaviour policy as a Gaussian Mixture Model (GMM). GMMs are flexible in representing complex distributions and can approximate any density function with sufficient components (McLachlan & Basford, 1988), making them well-suited for capturing the heterogeneous nature of offline datasets that often arise from multiple policies or exploration strategies. The behaviour policy $\pi_{b}(a\mid s)$ is expressed as:
@@ 4.2 Hyper-Markov Decision Process
$\mathcal{P}_{H}(s_{t+1}\mid s_{t},u_{t})=\mathbb{E}_{a_{t}\sim\phi^{u_{t}}(a\mid s_{t})}\left[\mathcal{P}(s_{t+1}\mid s_{t},a_{t})\right],$
$r_{H}(s_{t},u_{t})=\mathbb{E}_{a_{t}\sim\phi^{u_{t}}(a\mid s_{t})}\left[r(s_{t},a_{t})\right].$
@@ 4.3 Hyper Q-function
The proof can be found in Appendix B.1.
@@ 4.5 Weighted Imitation Learning on One Mode
Based on the greedy hyper-policy $\zeta_{g}$, we obtain the corresponding greedy policy $\pi_{\zeta_{g}}$, which, according to Theorem 1, is not worse than $\pi_{b}$. However, rather than simply imitating $\pi_{\zeta_{g}}$, we apply weighted imitation learning to further improve the policy. Specifically, we aim to find a policy $\pi$ that maximises the expected improvement $\eta(\pi)=J(\pi)-J(\pi_{\zeta_{g}})$. The expected improvement $\eta(\pi)$ can be expressed in terms of the advantage $A^{\pi_{\zeta_{g}}}(s,a)$ (Kakade & Langford, 2002; Schulman et al., 2015): $\eta(\pi)=\mathbb{E}_{s\sim d_{\pi}(\cdot),a\sim\pi(\cdot\mid s)}[A^{\pi_{\zeta_{g}}}(s,a)]$. However, in offline RL, state samples from the distribution $d_{\pi}(s)$ are unavailable. Instead, we approximate the expected improvement by using an estimate $\hat{\eta}$ (Kakade & Langford, 2002): $\hat{\eta}(\pi)=\mathbb{E}_{s\sim d_{\pi_{\zeta_{g}}}(\cdot),a\sim\pi(\cdot\mid s)}[A^{\pi_{\zeta_{g}}}(s,a)]$. This approximation provides a good estimate of $\eta(\pi)$ if $\pi$ and $\pi_{\zeta_{g}}$ are close in terms of KL-divergence (Schulman et al., 2015). As a result, we can formulate the following constrained policy optimisation problem:
$\begin{split}&\arg\max_{\pi}\sum_{s}d_{\pi_{\zeta_{g}}}(s)\sum_{a}\pi(a\mid s)A^{\pi_{\zeta_{g}}}(s,a),\\ &\text{s.t. }\sum_{s}d_{\pi_{\zeta_{g}}}(s)D_{KL}(\pi(\cdot\mid s)||\pi_{\zeta_{g}}(\cdot\mid s))\leq\epsilon.\end{split}$ (9)
@@
$\arg\min_{\pi}\mathbb{E}_{s\sim d_{\pi_{b}}(\cdot)}[D_{KL}(\pi^{*}(\cdot\mid s)||\pi(\cdot\mid s))]$ (10)
@@
In this framework, the actions to be imitated are sampled from $\pi_{\zeta_{g}}$, a unimodal Gaussian policy, thereby addressing the multi-modality issue. Furthermore, actions within this mode are weighted according to their advantage, encouraging the policy to concentrate on the highly-rewarded actions. As shown in Theorem 2, the policy learned by LOM has a theoretical advantage over both $\pi_{\zeta_{g}}$ and the original behaviour policy $\pi_{b}$.
@@
This theorem implies that our method introduces a novel two-step improvement process. LOM achieves further policy improvement compared to one-step algorithms while requiring less off-policy evaluation than multi-step algorithms (Brandfonbrener et al., 2021).